OpenStudy (anonymous):

Find the slope of every line that is perpendicular to the graph of the equation. y-5x=7

OpenStudy (texaschic101):

First we need to find the slope of the equation given by putting the equation in y = mx + b form, where m is your slope. Can you do that ?

OpenStudy (gorv):

first convert it in standard form y=mx+c y+5x=7 move 5x on other side then coefficient of x with its sign = value of m for two lines to be perpendicular product of their slope = -1 m* slope of perpendicular line = -1 slope of perpendicular line= -1/m